### v1

@@ 2.1 Problem Setup
Option $\text{Option Price}=\mathbb{E}_{\text{general risk neutral measure}}[f_{\text{payoff}}(\bm{Y})].$ (1)
Here $\bm{Y}$ represents a potentially high-dimensional random quantity that encapsulated all the randomness generated under a general, accessible sampling mechanism, denoted here by a general risk neutral measure $\mathbb{E}$. For example, if one is required to generate a discretized sample path of a certain stochastic differential equation (SDE) using some numerical scheme, $\bm{Y}$ then represents a complete sample path generated and the measure induced by that discretized SDE would be the general risk neutral measure. Typically, the more complex the simulation scheme is, the higher the dimension of $\bm{Y}$ will be, as more sophisticated simulation scheme often requires more intermediate steps or additional variables (e.g., an order of $O(T/\Delta t)$, when $T$ is the time horizon and $\Delta t$ is the discretization time step). Then, $f_{\text{payoff}}$, denoted by the term payoff function, represents the proper function (functional) that transform or summarize $\bm{Y}$ into a sample for the quantity of interest. The motivation for our framework arises, when the complexity and generality of the payoff function $f_{\text{payoff}}$ leads to a highly complicated dependence on the entirety of the random variables $\bm{Y}$. This complication may stem from factors such as path dependency, nested or multilevel simulations, or intricate interactions within the model. Consequently, drawing valid samples from the distribution of $f_{\text{payoff}}(\bm{Y})$ becomes both challenging and computationally intensive.
To fully specify the framework, we need to define the several parameters governing both the risk-neutral measure and the payoff function. We formally list the key definitions as:
  1. 1.
    Model Parameters $\bm{\theta}_{\text{model}}$: These are the parameters that specify the simulation or the stochastic model of the underlying process $\bm{Y}$ (e.g., asset value time series). For instance, in a Heston model (Heston 1993), the price process $S_{t}$ and the instantaneous volatility process $\nu_{t}$ are jointly modeled:
@@
    The model parameter $\bm{\theta}_{\text{model}}$ is thus 5 dimensional $\bm{\theta}_{\text{model}}:=(r,\eta,\delta,\rho,\kappa)\in\mathbb{R}^{5}$. It specifies the risk neutral drift $r$, the long-run average variance $\eta$, the volatility of volatility $\delta$, the correlation $\rho$ of between $W^{S}$ and $W^{\nu}$, and the mean reversion rate $\kappa$. In options pricing, these parameters are typically calibrated to fit market data and then used to generate sample paths $\bm{Y}:=\{S_{t}\}_{t}$. Thus $\bm{\theta}_{\text{model}}$ governs and describes $\mathbb{E}_{\text{ risk neutral measure}}$ that generates $\bm{Y}$.
  2. 2.
    Simulation Parameters $\bm{\theta}_{\text{simulation}}$: These parameters also specify $\mathbb{E}_{\text{risk neutral measure}}$. The distinction between $\bm{\theta}_{\text{simulation}}$ and $\bm{\theta}_{\text{model}}$ is somewhat artificial, primarily stemming from the convention that not all simulation parameters are not calibrated externally, but rather serve as hyperparameters of the Monte Carlo simulation itself (either customized or directly observed). For example, when one simulates the Heston model above using Euler discretization scheme, [^1] As mentioned in Dingec et al. (2015), Euler scheme was the primary method for simulation of Heston model until the exact simulation of Broadie and Kaya (2006) is developed. However, due to its significant computational demands, the exact method is less commonly used in practice and simpler, biased simulation method such as Quadratic Exponential (Andersen 2008) or Euler scheme remains prevalent (see, e.g., Lord et al. (2010)). we use $\bm{\theta}_{\text{simulation}}$ to specify the initial stock price $S_{0}$, the initial volatility $\nu_{0}$, the time horizon $T$, and discretization time step $\Delta t$. The choice of these parameters can impact both the accuracy and efficiency of the simulation. While not typically part of the model calibration process, these parameters also play a crucial role in describing risk-neutral measure.
  3. 3.
    Payoff Function Parameters $\bm{\theta_{\text{payoff}}}$: This parameter $\bm{\theta_{\text{payoff}}}$ parametrizes the payoff function $f_{\bm{\theta}_{\text{payoff}}}$. In options pricing, this specifies the details of the contract. For example, given the full price path $\bm{Y}:=\{S_{t}\}_{t}$, the payoff function for an arithmetic Asian options takes the form
@@
    where $\bm{\theta_{\text{payoff}}}$ specifies the strike level $K$, the sampling frequency $n_{D}$ and the observation dates $\{t_{i}\}_{i\in[n_{D}]}$. Notice we have simply omitted the discount factor $e^{-rT}$.
Under these definitions, we can summarize the standard Monte Carlo approach as follows: one calibrates $\bm{\theta}_{\text{model}}$, specify $\bm{\theta}_{\text{simulation}}$ and generate $n$ independent paths $\{\bm{Y}_{i}\}_{i\in[n]}$ according to $\mathbb{E}_{\text{ risk neutral measure}}$ given by $\bm{\theta}_{\text{model}}$ and $\bm{\theta}_{\text{simulation}}$. Finally, one averages $\frac{1}{n}\sum_{i}f_{\bm{\theta}_{\text{payoff}}}(\bm{Y}_{i})$.
@@
Remark Having specified the distinct roles of $\bm{\theta}_{\text{model}}$, $\bm{\theta}_{\text{simulation}}$ and $\bm{\theta}_{\text{payoff}}$, we will henceforth refer to them collectively as $\bm{\theta}$ when no ambiguity arises.
2.2 Prediction-Enhanced Monte Carlo: Motivation
As we have seen, Asian options are path-dependent derivative securities whose payoffs depend on the average prices of the underlying asset. Closed-form solutions for their prices are generally unavailable, numerous studies have proposed new numerical methods for pricing them (see Dingec et al. (2015)). While Partial Differential Equation (PDE)-based methods and Monte Carlo simulation are both generic methods for pricing options, applying PDE methods to particular models often requires special effort, primarily due to the curse of dimensionality. Moreover, for our discussion here, it has been noted that no PDE method performs well for Asian options under the Heston model Dingec et al. (2015). Therefore, across various models, simulation becomes the common choice. The principal drawback of simulation, as strongly highlighted in the literature, is its slow convergence rate. Employing variance reduction methods is therefore essential to improve the precision of simulation results. However, designing a successful variance reduction method requires exploiting the special structures of the derivative security and the underlying model. As a result, control variate (CV) techniques are often not available for pricing options under most of the complicated models used in practice. In fact, the pricing of arithmetic Asian option under Black–Scholes geometric Brownian motion (GBM) model is a main exception where the closed-form expression of the geometric Asian option allows one to effectively deploy control variate techniques. However, as soon as one extends beyond the most basic GBM model to more complex models such as the Heston model, the closed-form solutions for the geometric Asian options are no longer available and the applicability of control variate in these contexts is severely limited (see Dingec et al. (2015)).
More generally in the context of options pricing, while widely used, Monte Carlo methods face challenges when applied to complex, path-dependent derivatives. In such simulations, the state at each time step is determined by the state at the previous step, creating a sequential dependency that precludes straightforward parallelization or speedup. Moreover, for the complex risk neutral measures or the irregular payoff functions encountered in many exotic options pricing problems, effective control variate or variance reduction techniques are frequently unavailable. As a result, the slow convergence of Monte Carlo - its $O(1/\sqrt{n})$ error after $n$ replications - becomes an acute bottleneck and achieving high accuracy often implies a undesirable large number of replications in practice.
@@ 2.2 Prediction-Enhanced Monte Carlo: Motivation
To address these challenges, the Prediction-Enhanced Monte Carlo (PEMC) framework is designed to provide a robust and easily accessible mechanism for designing control variates. This novel approach aims to reduce the number of required Monte Carlo samples by generating a machine learning-based control variate that is both highly parallelizable and efficiently evaluable. Crucially, our framework preserves the unbiased nature of the Monte Carlo estimator, a property that Glasserman (2013b) highlights as “fundamental to the wide applicability of Monte Carlo methods in finance,” achieving a remarkable elimination of any potential bias introduced by any black-box prediction model.
2.3 Prediction-Enhanced Monte Carlo: Prediction Model
In this section, we present our PEMC framework, which is centered around an efficient prediction model. In PEMC, we do not specify the training algorithm or model architecture for this prediction model, as our framework is essentially compatible with any black-box model. However, as we will demonstrate, the squared error loss—a function that is both convex and differentiable—combined with the need for a flexible model family capable of expressing complex dependencies and being straightforward to train using off-the-shelf, easily accessible solvers, makes neural networks an ideal choice.
When applying the PEMC framework, an efficient prediction model refers to one that can process highly parallelizable and computationally efficient input $\bm{X}$, which serves as a strong predictor or good feature Hastie et al. (2009) for predicting $f_{\bm{\theta}_{\text{payoff}}}(\bm{Y})$. In particular, given the whole path $\bm{Y}$, we want to find some input feature $\bm{X}:=\phi(\bm{Y})$ that is some typically low-dimensional transformation $\phi$ of $\bm{Y}$ with:
  • Property 1: $\bm{X}$ can be regarded as a reasonable predictor for $f_{\bm{\theta}_{\text{payoff}}}(\bm{Y})$. For example, if $\bm{X}$ exhibits a significant correlation with $f_{\bm{\theta}_{\text{payoff}}}(\bm{Y})$, it may serve as the basis for a reasonable linear regressor, although our framework far exceeds linear predictors.
@@ 2.3 Prediction-Enhanced Monte Carlo: Prediction Model
  • Property 2: While the the joint distribution of $d$ is hard to obtain, the marginal distribution of $\bm{X}$ allows for efficient and parallelizable simulation.
The selection of an appropriate $\bm{X}$ is a pivotal component of our framework, analogous to feature selection in machine learning tasks. This process necessitates domain expertise and a thoughtful evaluation of the specific financial instrument and underlying model. The challenge lies in striking a balance between the predictive capability of $\bm{X}$ for $f_{\bm{\theta}_{\text{payoff}}}(\bm{Y})$ and its computational efficiency. In subsequent sections, we will explore the intricacies of this selection process in greater detail, examining its impact on the performance of the PEMC framework through illustrative examples. For now, we proceed under the assumption that a suitable $\bm{X}$ has been identified.
We now focus on the pre-training phase of our prediction model. Unlike many machine learning or deep learning tasks, where the quantity and quality of data are critical challenges—and often a core difficulty due to the expense or scarcity of labeled data—in the Monte Carlo setting, we benefit from an effectively unlimited supply of clean, labeled data that can be generated as needed. The data generation and model training process can be outlined through the following key steps:
  1. 1.
@@
    We first define a parameter space $\bm{\Theta}$ that encompasses all combinations of $\bm{\theta}:=(\bm{\theta}_{\text{model}},\bm{\theta}_{\text{simulation}},\bm{\theta}_{\text{payoff}})$ that are of practical interest. This space should cover the range of parameters that would likely be encountered in real-world pricing scenarios on a daily, weekly or monthly basis, depending on the model update frequency.
  2. 2.
@@
    We then draw samples of $\bm{\theta}_{i}$ from $\bm{\Theta}$, using uniform sampling or any measure that ensures comprehensive coverage. For each sampled parameter set, we generate one pair  [^2] You could also generate multiple pair of $(\bm{Y}(\bm{\theta}_{i}),\bm{X}(\bm{\theta}_{i}))$ under the same sampled $\bm{\theta}_{i}$. of $\bm{Y}(\bm{\theta}_{i})$ and its corresponding $\bm{X}(\bm{\theta}_{i}):=\phi(\bm{Y}(\bm{\theta}_{i}))$. It’s important to note that $\bm{X}$ and $\bm{Y}$ are coupled in this generation process.
  3. 3.
@@
    For each generated pair, we compute the payoff $f_{\bm{\theta}_{i,\text{payoff}}}(\bm{Y}(\bm{\theta}_{i}))$, which serves as a label in our training data. The corresponding feature vector is comprised of $(\bm{\theta}_{i,\text{model}},\bm{\theta}_{i,\text{simulation}},\bm{\theta}_{i,\text{payoff}},\bm{X}(\bm{\theta}_{i}))$.
  4. 4.
@@
    $\textit{label}_{i}:=$ $(\bm{\theta}_{i,\text{model}},\bm{\theta}_{i,\text{simulation}},\bm{\theta}_{i,\text{payoff}},\bm{X}(\bm{\theta}_{i}))$
    $\textit{feature}_{i}:=$ $f_{\bm{\theta}_{i,\text{payoff}}}(\bm{Y}(\bm{\theta}_{i}))$
@@
  5. 5.
    The machine learning model $g$ is then trained to minimize the mean squared error (MSE) loss and save the resulting model ${g}$:
@@
1:procedure DataGeneration($N_{\text{train}},\bm{\Theta}$)
@@
4:         Sample parameters $\bm{\theta}_{i}\sim\Theta$ uniformly
5:         Generate $\bm{X}(\bm{\theta}_{i}),\bm{Y}(\bm{\theta}_{i})\sim\mathbb{E}_{\text{ risk neutral measure}(\bm{\theta}_{i})}$
6:         $\textit{label}_{i}\leftarrow f_{\bm{\theta}_{i,\text{payoff}}}(\bm{Y}(\bm{\theta}_{i}))$
7:         $\textit{feature}_{i}\leftarrow(\bm{\theta}_{i,\text{model}},\bm{\theta}_{i,\text{simulation}},\bm{\theta}_{i,\text{payoff}},\bm{X}(\bm{\theta}_{i}))$
@@
The essential steps are summarized in Algorithm 1. While the data generation process can be time-intensive due to the complexity of $\bm{Y}$, it is conducted offline during the pre-training phase. This approach minimizes its impact on real-time pricing applications by effectively trading memory usage for computational speed during execution.
This pre-training setup presents an ideal scenario for machine learning applications. Unlike many real-world problems where data scarcity is a concern, access to the data-generating process in PEMC allows for the creation of virtually unlimited, clean data. This abundance greatly facilitates the training of complex models. It is important to also note that PEMC does not require $g$ to take specific forms. However, the convex and differentiable nature of the loss function (i.e., squared loss), the large supply of data (i.e., from Monte Carlo simulation), and the need for expressiveness of the model family make neural networks (NNs) an ideal choice.
When training NNs, various optimization techniques and architectural enhancements—such as the Adam optimizer Kingma and Ba (2014), batch normalization Ioffe and Szegedy (2015), and skip connections He et al. (2016)—can be employed to improve performance. In fact, recent advancements in machine learning technologies have made the training process remarkably accessible and efficient. The availability of open-source, user-friendly frameworks like PyTorch Paszke et al. (2019), TensorFlow Abadi et al. (2016), and JAX Bradbury et al. (2018)—coupled with GPU acceleration and highly optimized C++ backends—has dramatically simplified the process of training models on large datasets. These frameworks support recent neural network architectures, such as Convolutional Neural Networks (CNNs) LeCun et al. (1998), Residual Networks (ResNets) He et al. (2016), and Transformers Vaswani et al. (2017), as well as various Stochastic Gradient Descent (SGD) optimization algorithms Bottou (2010). In applications, we only utilized well-established and widely recognized neural network structures that are now considered common. These models can typically be implemented using just a few lines of code, as they have been incorporated into existing packages and extensively optimized.
@@
$g(\bm{\theta}_{\text{model}},\bm{\theta}_{\text{simulation}},\bm{\theta}_{\text{payoff}},\bm{X}(\bm{\theta}))$
$\approx$ $\mathbb{E}_{\text{risk neutral measure}(\bm{\theta})}\bigg{[}f_{\bm{\theta}_{\text{payoff}}}(\bm{Y}(\bm{\theta}))\bigg{|}(\bm{\theta}_{\text{model}},\bm{\theta}_{\text{simulation}},\bm{\theta}_{\text{payoff}},\bm{X}(\bm{\theta}))\bigg{]}$
@@
2.4 Prediction-Enhanced Monte Carlo: Evaluation

### v2

@@ 3.1. Problem Setup
Option $\text{Option Price}(\boldsymbol{\theta})=\mathbb{E}_{\text{risk neutral measure}(\boldsymbol{\theta})}[f_{\text{payoff}(\boldsymbol{\theta})}(\boldsymbol{Y})],$ (3)
For Asian option, we can specify the framework as:
  1. (1)
    Model Parameters $\boldsymbol{\theta}_{\text{model}}$: These are the parameters that specify the simulation or the stochastic model of the underlying process $\boldsymbol{Y}$ (e.g., asset value time series). For instance, in a Heston model [69], the price process $S_{t}$ and the instantaneous volatility process $\nu_{t}$ are jointly modeled:
@@
    The model parameter $\boldsymbol{\theta}_{\text{model}}$ is thus 5 dimensional $\boldsymbol{\theta}_{\text{model}}:=(r,\eta,\delta,\rho,\kappa)\in\mathbb{R}^{5}$. It specifies the risk neutral drift $r$, the long-run average variance $\eta$, the volatility of volatility $\delta$, the correlation $\rho$ of between $W^{S}$ and $W^{\nu}$, and the mean reversion rate $\kappa$. In options pricing, these parameters are typically calibrated to fit market data and then used to generate sample paths $\boldsymbol{Y}:=\{S_{t}\}_{t}$. Thus $\boldsymbol{\theta}_{\text{model}}$ governs and describes $\mathbb{E}_{\text{ risk neutral measure}}$ that generates $\boldsymbol{Y}$.
  2. (2)
    Simulation Parameters $\boldsymbol{\theta}_{\text{simulation}}$: These parameters also specify $\mathbb{E}_{\text{risk neutral measure}}$. The distinction between $\boldsymbol{\theta}_{\text{simulation}}$ and $\boldsymbol{\theta}_{\text{model}}$ is somewhat artificial, primarily stemming from the convention that not all simulation parameters are calibrated externally, but rather serve as hyperparameters of the Monte Carlo simulation itself (either customized or directly observed). For example, when one simulates the Heston model above using Euler discretization scheme, [^1] As mentioned in [44], Euler scheme was the primary method for simulation of Heston model until the exact simulation of [31] is developed. However, due to its significant computational demands, the exact method is less commonly used in practice and simpler, biased simulation method such as Quadratic Exponential [3] or Euler scheme remains prevalent (see, e.g., [91]). we use $\boldsymbol{\theta}_{\text{simulation}}$ to specify the initial stock price $S_{0}$, the initial volatility $\nu_{0}$, the time horizon $T$, and discretization time step $\Delta t$. The choice of these parameters can impact both the accuracy and efficiency of the simulation. While not typically part of the model calibration process, these parameters also play a crucial role in describing risk-neutral measure.
  3. (3)
    Payoff Function Parameters $\boldsymbol{\theta_{\text{payoff}}}$: This parameter $\boldsymbol{\theta_{\text{payoff}}}$ parametrizes the payoff function $f_{\boldsymbol{\theta}_{\text{payoff}}}$. In options pricing, this specifies the details of the contract. For example, given the full price path $\boldsymbol{Y}:=\{S_{t}\}_{t}$, the payoff function for an arithmetic Asian options takes the form
@@
    where $\boldsymbol{\theta_{\text{payoff}}}$ specifies the strike level $K$, the sampling frequency $n_{D}$ and the observation dates $\{t_{i}\}_{i\in[n_{D}]}$. Notice we have simply omitted the discount factor $e^{-rT}$. Note that the discrete sampling defied by the payoff does not necessarily determine the Euler discretization in time and in fact the latter is usually much more frequent for accuracy.
@@
Remark Having specified the distinct roles of $\boldsymbol{\theta}_{\text{model}}$, $\boldsymbol{\theta}_{\text{simulation}}$ and $\boldsymbol{\theta}_{\text{payoff}}$, we will henceforth refer to them collectively as $\boldsymbol{\theta}$ when no ambiguity arises.
3.2. Prediction-Enhanced Monte Carlo: Motivation
@@
As we have seen, Asian options are path-dependent derivatives whose payoffs depend on the average prices of the underlying asset. Generally, closed-form solutions for these options are unavailable, prompting numerous studies to propose various numerical methods for their pricing. In particular, the arithmetic Asian option under the geometric Brownian motion (GBM) model is one of the rare cases where a closed-form geometric Asian option price exists and can serve as a control variate. Beyond simple models like GBM, closed-form CV solutions are unavailable [44]. To address these challenges, the Prediction-Enhanced Monte Carlo (PEMC) framework aims to reduce the number of required Monte Carlo samples by generating a machine learning-based control variate that is both highly parallelizable and efficiently evaluable. Crucially, our framework preserves the unbiased nature of the Monte Carlo estimator, a property that [60] highlights as “fundamental to the wide applicability of Monte Carlo methods in finance,” achieving a remarkable elimination of any potential bias introduced by any black-box prediction model.
In PEMC, we do not specify the training algorithm or model architecture for this prediction model, as our framework is compatible with any black-box model. However, we want to find some input feature $\boldsymbol{X}:=\phi(\boldsymbol{Y})$ that is some typically low-dimensional transformation $\phi$ of $\boldsymbol{Y}$ with:
  • Property 1: $\boldsymbol{X}$ can be regarded as a reasonable predictor for $f_{\boldsymbol{\theta}_{\text{payoff}}}(\boldsymbol{Y})$.
@@
  • Property 2: In contrast to that of $\boldsymbol{Y}$, the marginal distribution of $\boldsymbol{X}$ allows for efficient and parallelizable simulation.
The selection of an appropriate $\boldsymbol{X}$ is a pivotal component of our framework, analogous to feature selection in machine learning tasks. This process necessitates domain expertise. For now, we proceed under the assumption that a suitable $\boldsymbol{X}$ has been identified.
3.3. Prediction-Enhanced Monte Carlo: Prediction Model
We now focus on the pre-training phase of our prediction model. The data generation and model training process can be outlined through the following key steps:
  1. (1)
@@ 3.3. Prediction-Enhanced Monte Carlo: Prediction Model
    We first define a parameter space $\boldsymbol{\Theta}$ that encompasses all combinations of $\boldsymbol{\theta}:=(\boldsymbol{\theta}_{\text{model}},\boldsymbol{\theta}_{\text{simulation}},\boldsymbol{\theta}_{\text{payoff}})$ that are of practical interest. This space covers the range of parameters that would likely be encountered in real-world pricing scenarios on a daily, weekly or monthly basis, depending on the model update frequency.
  2. (2)
@@
    We then draw samples of $\boldsymbol{\theta}_{i}$ from $\boldsymbol{\Theta}$, using uniform sampling or any measure that ensures comprehensive coverage. For each sampled parameter set, we generate one pair  [^2] You could also generate multiple pair of $(\boldsymbol{Y}(\boldsymbol{\theta}_{i}),\boldsymbol{X}(\boldsymbol{\theta}_{i}))$ under the same sampled $\boldsymbol{\theta}_{i}$. of $\boldsymbol{Y}(\boldsymbol{\theta}_{i})$ and its corresponding $\boldsymbol{X}(\boldsymbol{\theta}_{i}):=\phi(\boldsymbol{Y}(\boldsymbol{\theta}_{i}))$. It’s important to note that $\boldsymbol{X}$ and $\boldsymbol{Y}$ are coupled in this generation process.
  3. (3)
@@
    For each generated pair, we compute the payoff $f_{\boldsymbol{\theta}_{i,\text{payoff}}}(\boldsymbol{Y}(\boldsymbol{\theta}_{i}))$, which serves as a label in our training data. The corresponding feature vector is comprised of $(\boldsymbol{\theta}_{i,\text{model}},\boldsymbol{\theta}_{i,\text{simulation}},\boldsymbol{\theta}_{i,\text{payoff}},\boldsymbol{X}(\boldsymbol{\theta}_{i}))$.
  4. (4)
@@
    $\textit{feature}_{i}:=$ $(\boldsymbol{\theta}_{i,\text{model}},\boldsymbol{\theta}_{i,\text{simulation}},\boldsymbol{\theta}_{i,\text{payoff}},\boldsymbol{X}(\boldsymbol{\theta}_{i}))$
    $\textit{label}_{i}:=$ $f_{\boldsymbol{\theta}_{i,\text{payoff}}}(\boldsymbol{Y}(\boldsymbol{\theta}_{i}))$
@@
  5. (5)
    The machine learning model $g$ is then trained to minimize the mean squared error (MSE) loss and saved:
@@
1:procedure DataGeneration($N_{\text{train}},\boldsymbol{\Theta}$)
@@
4:         Sample parameters $\boldsymbol{\theta}_{i}\sim\Theta$ uniformly
5:         Generate $\boldsymbol{X}(\boldsymbol{\theta}_{i}),\boldsymbol{Y}(\boldsymbol{\theta}_{i})\sim\mathbb{E}_{\text{ risk neutral measure}(\boldsymbol{\theta}_{i})}$
6:         $\textit{label}_{i}\leftarrow f_{\boldsymbol{\theta}_{i,\text{payoff}}}(\boldsymbol{Y}(\boldsymbol{\theta}_{i}))$
7:         $\textit{feature}_{i}\leftarrow(\boldsymbol{\theta}_{i,\text{model}},\boldsymbol{\theta}_{i,\text{simulation}},\boldsymbol{\theta}_{i,\text{payoff}},\boldsymbol{X}(\boldsymbol{\theta}_{i}))$
@@
The essential steps are summarized in Algorithm 1. While the data generation process can be time-intensive due to the complexity of $\boldsymbol{Y}$, it is conducted offline during the pre-training phase. This approach minimizes its impact on real-time pricing applications by effectively trading memory usage for computational speed during execution.
This pre-training setup presents an ideal scenario for machine learning applications. It is important to also note that PEMC does not require $g$ to take specific forms. However, the convex and differentiable nature of the loss function (i.e., squared loss), the large supply of data (i.e., from Monte Carlo simulation), and the need for expressiveness of the model family make neural networks (NNs) an ideal choice. For training NNs, various optimization techniques and architectural enhancements—such as the Adam optimizer [80], batch normalization [75], and skip connections [64]—can be employed to improve performance. In fact, recent advancements in machine learning technologies have made the training process remarkably accessible and efficient. The availability of open-source, user-friendly frameworks like PyTorch [103], TensorFlow [1], and JAX [24]—coupled with GPU acceleration and highly optimized C++ backends—has dramatically simplified the process of training models on large datasets. These frameworks support recent neural network architectures, such as Convolutional Neural Networks (CNNs) [83], Residual Networks (ResNets) [64], and Transformers [128], as well as various Stochastic Gradient Descent (SGD) optimization algorithms [21]. In applications, we only utilized well-established and widely recognized neural network structures that are now considered common. These models can typically be implemented using just a few lines of code, as they have been incorporated into existing packages and extensively optimized.
@@
$g(\boldsymbol{\theta}_{\text{model}},\boldsymbol{\theta}_{\text{simulation}},\boldsymbol{\theta}_{\text{payoff}},\boldsymbol{X}(\boldsymbol{\theta}))$
$\approx$ $\mathbb{E}_{\text{risk neutral measure}(\boldsymbol{\theta})}\bigg{[}f_{\boldsymbol{\theta}_{\text{payoff}}}(\boldsymbol{Y}(\boldsymbol{\theta}))\bigg{|}(\boldsymbol{\theta}_{\text{model}},\boldsymbol{\theta}_{\text{simulation}},\boldsymbol{\theta}_{\text{payoff}},\boldsymbol{X}(\boldsymbol{\theta}))\bigg{]}$
@@
3.4. Prediction-Enhanced Monte Carlo: Evaluation